7.) Von Bora Corporation​ (VBC) is expected to pay a $ 3.00 dividend at the end of this year. If you expect​ VBC's dividend to grow by 6​% per year forever and​ VBC's equity cost of capital is 10​%, then the value of a share of VBS stock is closest​ to:

8.) NoGrowth Industries presently pays an annual dividend of $ 1.00 per share and it is expected that these dividend payments will continue indefinitely. If​ NoGrowth's equity cost of capital is 8​%, then the value of a share of​ NoGrowth's stock is closest​ to:
Answer #1

7) given D0 = $3

Equity cost (r) = 10%

Dividend growth (g)= 6%

Dividend for next year D1= D0 * (1+g)

= 3*(1+6%)

= 3.18

Value of share= D1/(r-g)

= 3.18/(10%-6%)

=$ 79.5

8) given D0 = $1

r= 8%

g= 0%

Apply same formula

Dividend for next year = D0 *(1+g)

= 1*(1+0%)

= 1

Value of share = D1/(r-g)

= 1/(8%-0%)

= $12.5